# Comparing Targeting Strategies for Maximizing Social Welfare with Limited Resources

### v1

@@ 4.6   Families of Welfare Functions
In order to simulate policymakers with varying preferences for who to treat, we compare risk-based targeting to treatment-effect based targeting on some general utility/social welfare functions that fall under the category of ’weighted power mean functions’ as described in [27]. The weighted power mean $M(\mathbf{u};\mathbf{w},p)$ for $\mathbf{u}\in\mathbb{R}^{d}_{+}$, $\mathbf{w}\in\Delta_{d-1}$, and $p\in\mathbb{R}\cup\{\pm\infty\}$ is defined as:
$M(\mathbf{u};\mathbf{w},p)=\begin{cases}\left(\sum_{i=1}^{d}w_{i}u_{i}^{p}\right)^{1/p}&p\neq 0\\ \prod_{i=1}^{d}u_{i}^{w_{i}}&p=0\end{cases}$
We choose this family of welfare functions because functions in this family satisfies a standard set of axiomatic properties [27]. The parameter $p$ determines the specific type of welfare function:
  • Utilitarian welfare ($p=1$): $M(\mathbf{u};\mathbf{w},p=1)=\sum_{i=1}^{d}w_{i}u_{i}$
  • Nash welfare ($p=0$): $M(\mathbf{u};\mathbf{w},p=0)=\prod_{i=1}^{d}u_{i}^{w_{i}}$
We consider utilitarian welfare with two sets of weights. First, the uniform weights $w_{i}=1\forall i\in[n]$. Second, $w_{i}=\frac{n\cdot e^{\alpha b^{\prime}(X_{i})}}{\sum_{j=1}^{n}e^{\alpha b^{\prime}(X_{j})}}$ where $\alpha$ is a hyperparameter and $b^{\prime}(X_{i})$ is represents the percentile score of the baseline risk for the $i^{\text{th}}$ example, with the example with highest baseline risk having score 1 and the example with lowest baseline risk having score 0. This assigns greater weight to individuals with high values of baseline risk for high values of $\alpha$, thereby simulating a scenario where a policymaker might value treating treating these "high risk" individuals for other reasons. $\alpha$ can be interpreted as $2\log\left(\frac{w_{75}}{w_{25}}\right)$ where $w_{75}$ is the weight given to the $75^{\text{th}}$ percentile example by baseline risk and $w_{25}$ is the weight given to the $25^{\text{th}}$ percentile example by baseline risk.
The Nash social welfare function has commonly been used to achieve a balance between maximizing total welfare(utilitarian) and ensuring equitable distribution(egalitarian) [6, 7]. Egalitarian welfare can sometimes leads to inefficiencies while utilitarian welfare can lead to unjust outcomes: Nash welfare often strikes a useful compromise between these two ends. We consider unweighted Nash welfare ($w_{i}=1\forall i$). In order to avoid the complications of utility when using an unweighted Nash welfare function, we scale up the estimated utilities for each example to a minimum value of 1.
Note that the Nash welfare can be equivalently formulated in log space [6] as
$\frac{1}{n}\sum_{i=1}^{n}\log u_{i}.$
When each individual’s utility under an allocation policy corresponds to their realized outcome $Y_{i}^{(A(X_{i}))}$ (e.g., their income after the interventional period), we compare policies exactly as outlined in Equation 4.5, but with $Y^{(A(X))}$ replaced by $\log Y^{(A(X))}$, estimated by replicating the same procedure after taking logs of all outcome variables.
@@ 5 Results
  • The Ultra Poor Dataset exhibits highest treatment effects at high levels of baseline risk: treatment effect first decreases then increases consistently before slightly tapering off.
  • The NSW Dataset does not show any apparent trends in treatment effect as a function of baseline risk with the curve being mostly flat.
  • The Postoperative Pain Dataset shows a sigmoid-like curve, with treatment effects increasing more rapidly at moderate levels of baseline risk.
  • For the Acupuncture Dataset, we see treatment effects peaking at moderate levels of baseline risk after an initial dip and then slightly decreasing again before taking on mostly similar values for examples with high baseline risk.
  • The Tennessee STAR Dataset displays a generally increasing trend in treatment effects with baseline risk, but with notable fluctuations.
These diverse patterns underscore the complexity of the relationship between baseline risk and treatment effectiveness. Rather than assuming a particular relationship, it is crucial to empirically investigate the heterogeneity of treatment effects for each specific intervention and population.
@@
Figure 2 shows the comparison between risk-based targeting and treatment effect based targeting for the 5 datasets we study, with varying degrees of confounding and under different social welfare functions used to represent policymaker preferences. For reference, we also show the performance of a random targeting strategy which allocates the same budget but targets uniformly random individuals. Both targeting methods seem to perform better than random targeting for almost all combinations of dataset/welfare functions we consider. The improved performance of risk-based targeting over random reflects the generally increasing relationship that Figure 1 shows between baseline risk and treatment effects. However, when treatment effects can be accurately estimated ($k=0$, no confounding), targeting based on treatment effects always produces significantly higher utilitarian welfare (often producing a treatment-on-the-treated effect of three times or more greater than risk-based targeting). This indicates that when a policymaker seeks only to maximize aggregate benefit and can credibly estimate treatment effects, the gains from causal targeting are substantial.
@@
The second column of Figure 2 shows an alternative set of preferences, where the policymaker has a weighted utilitarian welfare function (for these plots, $\alpha$ is $2\log 2$, the value at which the ratio of the 75th percentile weight to the 25th percentile weight is 2) which places greater weight on individuals with higher baseline risk, attaching a higher importance to the welfare of more vulnerable individuals. This welfare function decreases the gap between treatment effect and risk based targeting as individuals with higher risk now have more weight associated with them. However, targeting on (biased) treatment effects is still preferable to risk-based targeting across all datasets.
@@
Table I: Values of $\alpha$ for different k at which risk-based targeting outperforms treatment effect based targeting. ’na’ indicates no such $\alpha$ was found
@@
Under the Nash social welfare function (third column of Figure 2), we observe some differences across our chosen settings but the general trend remains the same: as we increase confounding (increase the percentage of data we systematically remove), the benefit we accrue by following a treatment assignment policy based on biased treatment effect values decreases. At high levels of confounding, risk-based targeting accrues higher utility for a policymakers with a Nash social welfare function in the Ultra Poor setting in 2. However, across all other datasets, the policymaker prefers to target based on treatment effects even at high levels of confounding.
Collectively, these results indicate that policymakers are almost always better off targeting interventions based on treatment effect estimates rather than baseline risk values from a machine learning model, unless they are extremely inequality-averse.
@@ 6 Conclusion
This paper presents a systematic comparison between two of the most popular treatment assignment policies in use by policymakers today in risk-based targeting and treatment effect based targeting. Using data collected in 5 randomized controlled trials in different settings, we propose a systematic method of introducing confounding in our data and show that policies that rely on biased estimates of treatment effect from confounded data gradually perform worse as we further confound the data. We also provide empirical evidence of non-monotonic heterogeneity in treatment effect as a function of the baseline risk, implying that risk-based targeting is often suboptimal from the perspective of maximizing the overall effect of treatment assignment. Importantly, our analysis extends beyond a simple unweighted utilitarian approach, encompassing weighted utilitarian and Nash welfare functions that respectively prioritize high-risk individuals and fairness. We argue that while risk-based targeting can be effective in certain contexts, especially when RCTs are infeasible and the observational data available is heavily confounded, incorporating even biased information about treatment effects is likely to result in improved welfare.
@@ Appendix A Appendix
A.1   Datasets
@@
A.2   Additional Plots

### v2

@@ 3.5   Families of Welfare Functions
In order to simulate policymakers with varying preferences for who to treat, we compare risk-based targeting to treatment-effect based targeting with respect to two paradigmatic classes of social welfare functions. A social welfare function maps a vector of individual utilities $\mathbf{u}$ to the policymaker’s overall utility for the allocation. Both functions we study are commonly used instances of the class of weighted power mean functions, which contains all social welfare functions satisfying desirable axiomatic properties [30]. The two functions we investigate here are Weighted Utilitarian Welfare, given by $M(\mathbf{u};\mathbf{w})=\sum_{i=1}^{d}w_{i}u_{i}$ and Nash Welfare, given by $M(\mathbf{u})=\prod_{i=1}^{d}u_{i}$.
We consider utilitarian welfare with two sets of weights. First, the uniform weights $w_{i}=1\forall i\in[n]$. Second, $w_{i}=\frac{n\cdot e^{\alpha b^{\prime}(X_{i})}}{\sum_{j=1}^{n}e^{\alpha b^{\prime}(X_{j})}}$ where $\alpha$ is a hyperparameter and $b^{\prime}(X_{i})$ is represents the percentile score of the baseline risk for the $i^{\text{th}}$ example, with the example with highest baseline risk having score 1 and the example with lowest baseline risk having score 0. This assigns greater weight to individuals with high values of baseline risk for high values of $\alpha$, thereby simulating a scenario where a policymaker might value treating treating these "high risk" individuals. $\alpha$ can be interpreted as $2\log\left(\frac{w_{75}}{w_{25}}\right)$ where $w_{75}$ is the weight given to the $75^{\text{th}}$ percentile example by baseline risk and $w_{25}$ is the weight given to the $25^{\text{th}}$ percentile example by baseline risk.
The Nash social welfare function has commonly been used to achieve a balance between maximizing total welfare (utilitarian) and ensuring equitable distribution (egalitarian) [6, 7]. We consider unweighted Nash welfare ($w_{i}=1\forall i$). In order to avoid the complications of utility when using an unweighted Nash welfare function, we scale up the estimated utilities for each example to a minimum value of 1. Note that the Nash welfare can be equivalently formulated in log space [6] as $\frac{1}{n}\sum_{i=1}^{n}\log u_{i}$. When each individual’s utility under an allocation policy corresponds to their realized outcome $Y_{i}^{(A(X_{i}))}$ (e.g., their income after the interventional period), we compare policies exactly as outlined in Equation 2, but with $Y^{(A(X))}$ replaced by $\log Y^{(A(X))}$, estimated by replicating the same procedure after taking logs of all outcome variables.
@@ 4 Results
Next, Figure 2 shows the the average utility of the three targeting policies – risk-based, treatment effect-based, and random – on each of the datasets. The $x$ axis on each plot represents the degree of confounding synthetically introduced, as discussed above. At the $k=0$ point ($x=0$) on the axis, no confounding is present, representing the ideal scenario as in an RCT. For risk-based targeting, we observe a tendency towards improvement over random targeting, but with confidence intervals overlapping in all domains. For treatment-effect based targeting, we observe at most mixed success. On the STAR dataset, we see statistically significant evidence that targeting on treatment effects outperforms risk-based targeting when no confounding is present, with the point estimate trending downwards and eventually drawing even with risk-based targeting as confounding increases. On the NSW dataset, we see a similar trend in the point estimates, but with overlapping confidence intervals. One the remaining datasets, treatment effect targeting has strongly overlapping confidence intervals with risk-based targeting and shows effectively a flat trajectory with respect to the degree of confounding. We conclude that in these settings, either limited data or insufficiently rich features prevent the model from learning complex treatment effect heterogeneity patterns. On the one hand, we conclude that practitioners should be wary about whether the kind of data generated by many RCTs in practice suffices to learn strong predictors of heterogeneous treatment effects. On the other hand, we find no evidence that the presence of confounding makes matters worse unless there is a meaningful amount of signal present to start with. Appendix A contains results for alternate welfare functions, with similar results.
@@
Figure 3: Comparison of risk-based targeting to biased treatment effect-based targeting by plotting the benefit offered by each policy against the amount of data systematically removed from the RCT to introduce confounding but using observed pseudo-outcome data
In order to isolate the impact of confounding in a setting where heterogeneous treatment effects would otherwise be predictable, we now conduct a semi-synthetic analysis that simulates a targeting model with stronger signal. In particular, we simulate a treatment effect-based targeting policy that ranks individuals by their pseudo-outcome directly, instead of the predicted pseudo-outcome using the covariates. This incorporates information on the realized outcome of the individual and so is not attainable in practice. However, it mimics a setting where the second-stage regression model has strong performance in the no-confounding setting. Differences between the two settings are detailed in A.1. The results, shown in Figure 3, demonstrate that when treatment effects can be accurately estimated ($k=0$, no confounding), targeting based on treatment effects always produces significantly higher utilitarian welfare. This indicates that when a policymaker seeks only to maximize aggregate benefit and can credibly estimate treatment effects, the gains from causal targeting are substantial.
@@
The second column of Figure 3 shows an alternative set of preferences, where the policymaker has a utilitarian welfare function with higher welfare weights on individuals with higher baseline risk (for these plots, $\alpha$ is $2\log 2$, so the ratio of the weight placed on individuals in the 75th percentile of risk to the 25th percentile is 2). This welfare function decreases the gap between treatment effect and risk based targeting as individuals with higher risk now have more weight associated with them. However, targeting on (biased) treatment effects is still preferable to risk-based targeting across all datasets in Figure 3. Similarly, under the Nash social welfare function (third column of Figure 3), we observe some differences across our chosen settings but the general trend remains the same: as we increase confounding (increase the percentage of data we systematically remove), the benefit we accrue by following a treatment assignment policy based on biased treatment effect values decreases. At high levels of confounding, risk-based targeting accrues higher utility for a policymakers with a Nash social welfare function in the Ultra Poor setting in 3. However, across all other datasets, the policymaker prefers to target based on treatment effects even at high levels of confounding.
This motivates us to ask: how much greater must the policymaker weight high-risk individuals in order to prefer risk-based targeting? In Table I we give the minimum value of $\alpha$ at which risk-based targeting finally outperformed treatment effect based-targeting at each level of systematic data removal for the real and synthetic settings we consider respectively. We limit $\alpha$ such that the ratio of the $75^{\text{th}}$ percentile weight to the $25^{\text{th}}$ percentile weight when sorted in descending order is less than $100$; otherwise, only a few individuals have non-zero weights and estimating welfare gains becomes impossible. “na" indicates there is no $\alpha$ value (up to our upper bound) at which risk based targeting outperforms targeting based on treatment effects.
@@
Table I: Values of $\alpha$ for different k at which risk-based targeting outperforms treatment effect based targeting. ’na’ indicates no such $\alpha$ was found (using observed pseudo-outcome data)
@@
Table II: Values of $\alpha$ for different k at which risk-based targeting outperforms treatment effect based targeting. ’na’ indicates no such $\alpha$ was found
Dataset 5% 10% 15% 20% 25% 30% 35% 40%
Ultra Poor 0 0 0 0 0 0 0 0
NSW na na na na 0 0 na na
Postoperative Pain 0 0 0 0 0 0 0 0
Acupuncture 3.5 3 na 9 6 7 7 3.5
STAR 5 3 3.5 2 3.5 2 0.5 1.5
Collectively, our results reveal two key insights. First, in many real-world settings, limited data makes it difficult to learn reliable mappings from features to treatment effects, which can severely limit the effectiveness of treatment effect based targeting in practice. However, our analysis using directly observed pseudo outcomes, as well as real-data results in the settings with stronger signal, both demonstrate that simple reliability of learning heterogeneous treatment effects is the key bottleneck: if treatment effects could be estimated more reliably (through larger datasets or better modeling approaches), then a policy maker is almost always better of using them for targeting, regardless of both the potential for confounding and inequality-averse preferences.
@@ 5 Conclusion
This paper presents a systematic comparison between two of the most popular treatment assignment policies in use by policymakers today: risk-based targeting and treatment effect based targeting. We find that risk-based targeting tends to produce higher welfare than a uniformly random allocation, confirming some of the intuition behind its widespread use by practitioners. However, our analysis reveals both important practical limitations and significant potential benefits of treatment effect based targeting. The key practical challenge is that real-world datasets can contain insufficient data to learn reliable mappings from features to treatment effects, making it difficult to accurately predict who will benefit most from treatment. This limitation helps explain why risk-based targeting, which requires learning simpler relationships, remains popular in practice.
At the same time, our analysis which simulates a higher-accuracy treatment effect estimator shows the substantial potential benefits of treatment effect based targeting if these estimation challenges could be overcome. Even when accounting for confounding in treatment effect estimates and egalitarian preferences for assisting high-risk individuals, treatment effect based targeting consistently produces better outcomes unless policymakers have extremely strong preferences for helping higher-risk individuals regardless of benefit. One limitation is that our investigation assumes an essentially consequentialist perspective, where the policymaker’s goal is to improve individuals’ welfare as defined by their outcome. If policymakers have non-consequentialist preferences, targeting directly on a measure of vulnerability may be more appropriate.
Looking forward, we nevertheless find that there would be substantial gains to investing in data and modeling approaches that enabled effective causal targeting (at least with respect to the normative goals we assess). Our results also suggest that enabling strong outcome prediction is worth accepting some risk of bias, potentially arguing in favor of increased use of observational administrative data.
Acknowledgement:
This work was supported in part by the AI2050 program at Schmidt Futures (Grant G-22-64474).
@@ Appendix A Appendix
A.1   Experiment Details
Real Setting:
We divide the RCT data into two splits such that one split is used for training nuisance functions and the other split is used entirely for evaluation.
Following the DR estimator framework in [20], we divide the first split into two halves and and train a set of outcome and propensity nuisance functions on each half. We then perform cross-fitting where the nuisance functions trained on one half are used to estimate pseudo outcomes for the other half. Finally, we train a random forest regressor strictly between the estimated pseudo outcome differences and a unit’s features. This random forest model is applied to units in the evaluation set to obtain treatment effect estimates for the evaluation set. The ground truth treatment effects for the evaluation set are obtained by using information about the observed pseudo outcome and taking the mean of the estimate provided by each set of nuisance functions.
Semi-synthetic Setting:
We divide the RCT into two splits such that we use each split to obtain treatment effect estimates for the other split and make maximal use of available data. Again, we use the DR estimator from [20] and use information about the observed pseudo outcome in estimating treatment effects (both ground truth and subsequent biased treatment effect estimates from confounded data).
A.2   Datasets
@@
A.3   Kernel Smoothing
$\sigma_{i}=\frac{1}{2}(b(X_{i+100})-b(X_{i-99}))$ (4)
$\text{CI}=\hat{\tau}(b(X_{i}))\pm 1.96\sqrt{\frac{\sum_{j=1}^{n}K(\frac{b(X_{j})-b(X_{i})}{\sigma_{i}})(\hat{\tau}_{j}-\hat{\tau}(b(X_{i})))^{2}}{(\sum_{j=1}^{n}K(\frac{b(X_{j})-b(X_{i})}{\sigma_{i}}))^{2}}}$ (5)
A.4   Additional Plots
A.4.1 Evaluating based on Weighted Utilitarian Welfare and Nash Social Welfare
Figure 4: Comparing risk-based targeting to biased treatment effect-based targeting on weighted utilitarian welfare and Nash social welfare for the STAR RCT.
Figure 5: Comparing risk-based targeting to biased treatment effect-based targeting on weighted utilitarian welfare and Nash social welfare for the Ultra Poor RCT.
Figure 6: Comparing risk-based targeting to biased treatment effect-based targeting on weighted utilitarian welfare and Nash social welfare for the NSW RCT.
Figure 7: Comparing risk-based targeting to biased treatment effect-based targeting on weighted utilitarian welfare and Nash social welfare for the Postoperative Pain RCT.
Figure 8: Comparing risk-based targeting to biased treatment effect-based targeting on weighted utilitarian welfare and Nash social welfare for the Acupuncture RCT.
[…]